OpenStudy (anonymous):

Dilbert's TV and Appliance discounts all televisions 7% to customers paying in cash. Rebecca paid $1,748.22 in cash and before tax for a 52-inch widescreen TV. What was the original price of the TV, not including tax?

OpenStudy (anonymous):

idk srry

OpenStudy (anonymous):

x = original price x = 1748.22 - .07x --- add .07x to both sides .07x + x = 1748.22 1.07x = 1748.22 x = 1748.22 / 1.07 x = 1633.85 check.. 1633.85 + .07(1633.85) = 1748.22 1633.85 + 114.37 = 1748.22 1748.22 = 1748.22 (correct) original price (x) was $ 1633.85